OpenStudy (anonymous):

Match each expression with its value. sin 5π/6

OpenStudy (anonymous):

I got 1/2

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

@nickhouraney last one. Is this right cos 2π/3= -1/2

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

if you in trig its best if you memorize the unit circle. theres a pattern so you dont have to literally memorize everything http://en.wikipedia.org/wiki/File:Unit_circle_angles_color.svg

OpenStudy (anonymous):

@nickhouraney sorry to bother you. I know to do the sin and cos but don't know how to do tan. tan 5π/3

OpenStudy (anonymous):

alryte so check out that link i sent you. sin of 5pie/3 = -rad3/2 , cos of 5pie/3 = 1/2. so all we have to do is sin/cos

OpenStudy (anonymous):

\[\frac{ \frac{ -\sqrt{3} }{ 2 } }{ \frac{ 1 }{ 2 } }\]

OpenStudy (anonymous):

then it will equal \[\frac{ -\sqrt{3} }{ 2 }\times \frac{ 2 }{ 2 }\]

OpenStudy (anonymous):

thats supposed to be a 2/1 on the right

OpenStudy (anonymous):

got it so far?